The light Ray is incident at an angle of60 on a prism of angle 45 when the light Ray fall s on the other surface at 90 the refractive index of the material of the prism new and angle of deviation d are given by 

Grade:12

1 Answers

Vikas TU
4 years ago
Hiii 
Here r1 = 45 degree
so, Sin i/Sin r1 = Sin 60 / sin 45
so, new = sqrt (3/2)
delta = i + e – A
delta = 60 + 0 – 45 = 15 degree